OpenStudy (anonymous):

please help me :) A Ferris wheel at a carnival has a diameter of 62 feet. Suppose it turns at a rate of 70 revolutions per hour. (a) Find the angular speed of the wheel in radians per hour. (b) Find the linear speed of a passenger in feet per minute. Do not round any intermediate computations, and round your answer to the nearest whole number.

OpenStudy (anonymous):

a- angler speed = 70 *2pie = 140*22/7 = 20*22 = 440 rad/hr

OpenStudy (anonymous):

ok so how do i do b?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

each round circumference = pie * diameter = 22/7 * 62 = 195 ft speed = 70 round/hr = 70 round / 60 minute = 7/6 round/ min so speed = (7/6)*195 feet/min = 227 feet/min

OpenStudy (anonymous):

yes!!! thank you, thank you!!! can you help me on another problem?

OpenStudy (anonymous):

yes you r welcome

OpenStudy (anonymous):

A ceiling fan has 19-inch blades (so the radius of the circular fan is 19 inches). Suppose the fan turns at a rate of 55 revolutions per minute. (a) Find the angular speed of the fan in radians per minute. (b) Find the linear speed of the tip of a blade in inches per minute.

OpenStudy (anonymous):

a - anguler speed = pie *55 = (22/7)*55 = 173 rad/minute

OpenStudy (anonymous):

b- each round = pie*diameter = (22/7)*19 = 60 inch speed = 55 round/minute = 55*60 inch/minute = 3300 inch/minute

OpenStudy (anonymous):

other questions

OpenStudy (anonymous):

yes but wait sorry i forgot to post my last msg i had to go to class.

OpenStudy (anonymous):

its says its wrong :(

OpenStudy (anonymous):

really

OpenStudy (anonymous):

which one is wrong

OpenStudy (anonymous):

really

OpenStudy (anonymous):

the previous question is right or not ?

OpenStudy (anonymous):

yea the previous one was right!

OpenStudy (anonymous):

yes yes yes sure it will wrong i use the diameter scale ...where question gave us radius

OpenStudy (anonymous):

oh see i have no idea of what I'm doing…

OpenStudy (anonymous):

ok this is the correct answer

OpenStudy (anonymous):

a- 346 b- 6600

OpenStudy (anonymous):

I'm seriously stressing right now I'm in this online math class and it gives me four question if i get them right i exit and succeed the but if i answer three out of four it makes me answer four more until there all right.

OpenStudy (anonymous):

its says b is wrong

OpenStudy (anonymous):

try 6569

OpenStudy (anonymous):

what is the result

OpenStudy (anonymous):

i only get one more chance to answer correctly...

OpenStudy (anonymous):

ok give me the question again

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it was right thanks

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is this exame fele" where if your name "

OpenStudy (anonymous):

i understand what you doing but i can't figure out how to do it on my own..

OpenStudy (anonymous):

what is your class or your education level

OpenStudy (anonymous):

ugh next question…. A cyclist is riding a bicycle whose wheels have a radius of 11 inches. Suppose she is traveling at 25 miles per hour. (A useful fact: =1mi5280ft.) (a) Find the angular speed of the wheels in radians per minute. (b) Find the number of revolutions the wheels make per minute.

OpenStudy (anonymous):

well it's trig but for some reason its asking physics questions.

OpenStudy (anonymous):

a-2405 rad/min

OpenStudy (anonymous):

b-383

OpenStudy (anonymous):

right or not

OpenStudy (anonymous):

its say one of the digits in the answers is wrong on both answers…

OpenStudy (anonymous):

so i tried working it out on my own and i got 1584000, and 26400

OpenStudy (anonymous):

1- 2201 2-350

OpenStudy (anonymous):

no…thats wrong.

OpenStudy (anonymous):

do you know how much exactly inch to feet equal to?

OpenStudy (anonymous):

foot --- inch

OpenStudy (anonymous):

well its said (one of the digits) not the whole prob.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there is problem in change ft to inch do you know how much

OpenStudy (anonymous):

1-2401 2-382

OpenStudy (anonymous):

so we got that wrong the answer ***** a. 2400 b.382*****

OpenStudy (anonymous):

thank you…. Suppose that Saturn rotates on its axis once every 10.7hours. The equator lies on a circle with a radius of 37,450miles. (a) Find the angular speed of a point on its equator in radians per day (24 hours). (b) Find the linear speed of a point on the equator in miles per day.

OpenStudy (anonymous):

im sorry for last wrong answer

OpenStudy (anonymous):

no no thats fine…. thank you so much for helping. my teacher couldn't even help me I'm just at a stand still until i complete this section.

OpenStudy (anonymous):

1-14 rad /day

OpenStudy (anonymous):

no no no wait

OpenStudy (anonymous):

1-235400 2-9808

OpenStudy (anonymous):

what is it... right ???????

OpenStudy (anonymous):

no the answer was a:14rad b:527,788

OpenStudy (anonymous):

like i give it to you first

OpenStudy (anonymous):

yes but i didn't see that so it add 4 more questions.

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

what is the qu.

OpenStudy (anonymous):

its fine you've helped me a lot i just have to test out…

OpenStudy (anonymous):

A Ferris wheel at a carnival has a diameter of 60feet. Suppose a passenger is traveling at 5 miles per hour. (A useful fact: =1mi5280ft.) (a) Find the angular speed of the wheel in radians per minute. (b) Find the number of revolutions the wheel makes per hour. (Assume the wheel does not stop.)